# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

### v1

@@ VII-B Impact of Sensing Requirement
Figure 7: Blocklength vs. sensing SINR threshold.
Figure 8: E2E energy consumption vs. sensing SINR threshold.
The impact of the sensing SINR threshold, $\gamma_{s}$, on the blocklength and E2E energy consumption is investigated in Fig. 7 and Fig. 8, respectively. E2E energy minimization leads to lower energy consumption as it is shown that E2E SeURLLC+ outperforms SeURLLC+. The reason can be found in Fig. 7. It is shown that E2E SeURLLC+ leads to lower blocklength which means lower E2E energy consumption. Moreover, we observe that the gap between the performance of these two algorithms gradually decreases by demanding higher sensing SINR threshold.
@@ VIII Conclusion
In this paper, we proposed a joint blocklength and power control algorithm for a downlink cell-free massive MIMO system with multi-static sensing and URLLC users. We formulated two non-convex optimization problems, focusing on transmission energy minimization and E2E energy minimization, respectively. Numerical results demonstrated the superiority of the E2E energy minimization algorithm, particularly in scenarios with no sensing capability. Furthermore, our results revealed a growing significance of energy consumption related to sensing processing tasks with the increase in the number of sensing receive access points. This observation highlights a potential optimization avenue, as deactivating one sensing receive AP can effectively mitigate this dominance and contribute to overall energy efficiency. Furthermore, our findings indicated that a higher sensing SINR threshold leads to increased E2E energy consumption. This ultimately narrows the performance gap between the two proposed algorithms: the transmission-only approach and the E2E energy minimization approach.
@@ Appendix A Proof of Theorem 1
Let us define a new optimization variable, denoted by $\overline{L}$, where $L-L_{p}\leq 1/\overline{L}$. Then, minimizing the objective function is equivalent to minimizing the following convex function (quadratic-over-linear function)
$\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}\,.$ (73)
Since the above function is minimized, at the optimal solution, it holds that $\overline{L}=1/L_{d}=1/(L-L_{p})$. According to (26), the reliability constraints in (63b) can be written in the form of
@@
where $\overline{\mathsf{SINR}}^{\rm(dl)}_{i}$ is substituted by (23). To handle the non-convexity of the left-hand side in (74), we define a new variable $\chi_{i}$ and use fractional programming [36] to write the left-hand side as
@@
which can be written as a second-order cone (SOC) constraint in (64k). We then re-cast the constraint in (75) as
@@
$\boldsymbol{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\boldsymbol{\rho}\!\leq\!-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}.$ (78)
Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (63e) in SOC form in terms of $\boldsymbol{\rho}$ as $\left\|\textbf{F}_{k}\boldsymbol{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ , for $k=1,\ldots,N_{\rm tx}$.

### v2

@@ VII-B Impact of Sensing Requirement
The impact of the sensing SINR threshold, $\gamma_{s}$, on the performance of the E2E SeURLLC+ algorithm is evaluated in Figs. 4a-c in terms of E2E energy consumption, detection probability, and sensing update rate, respectively. Comparing Fig. 4(a) with Fig. 4(b), we observe that both detection probability and energy consumption generally increase with a higher SINR threshold. The clutter-aware detector with $N_{\rm rx}=2$ provides the highest detection probability, consistently above 0.9, whereas the clutter-unaware detector cannot achieve a detection probability over 0.6, despite offering the lowest energy consumption. Notably, using a clutter-aware detector while disabling one RX-AP maintains a detection probability above 0.8 with around 43% energy saving. This suggests a potential optimization strategy: turning off one sensing receive AP can halve the energy consumption, significantly enhancing overall energy efficiency. Higher sensing SINR threshold demands higher blocklength, resulting in lower sensing update rate as shown in Fig. 4(c). Comparing Fig. 4(b) and Fig. 4(c), it is evident that increasing the sensing SINR threshold significantly decreases the sensing update rate without substantially improving detection probability.
Figure 4: (a) E2E energy consumption, (b) detection probability, and (c) sensing update rate vs. sensing SINR threshold with E2E SeURLLC+ algorithm.
VII-C Impact of Number of APs
We explore the impact of the number of transmit and receive APs on the network performance in terms of E2E energy consumption, detection probability, and sensing update rate as shown in Figs. 5a-c, respectively. The results in Fig. 5(a) and Fig. 5(b) indicate that increasing the number of receive APs generally raises both E2E energy consumption and detection probability. For $N_{\rm rx}=2$, the detection probability substantially improves with only a marginal raise in energy consumption. However, further increasing to $N_{\rm rx}=4$ offers little improvement in detection probability while significantly increasing energy consumption. Moreover, the results in Fig. 5(c) show a consistent sensing update rate across all curves, especially for $N_{\rm tx}=8$ and $N_{\rm tx}=12$, indicating that $N_{\rm rx}$ has minimal impact on $R_{s}$. This suggests the optimization algorithm is more constrained by URLLC requirements than by sensing requirements.
A comparison between $N_{\rm tx}=8$ and $N_{\rm tx}=12$ in Fig. 5(c) reveals that a higher $N_{\rm tx}$ enhances URLLC satisfaction due to elevated SINR, resulting in a reduced blocklength and consequently a higher sensing update rate. It is observed that when $N_{\rm rx}=1$ or $N_{\rm rx}=2$, increasing the number of transmit APs does not significantly affect E2E energy consumption. This is because the rise in energy consumption from the additional APs is offset by the decrease in blocklength. Thus, the overall E2E energy consumption remains relatively unchanged. Conversely, when $N_{\rm rx}=4$, the influence of $N_{\rm tx}$ becomes more pronounced, as the energy consumption for sensing tasks increases non-linearly with the product of $N_{\rm tx}$ and $N_{\rm rx}$.
When $N_{\rm tx}=16$, the system’s behavior concerning E2E energy consumption and detection probability follows a similar trend, showing an increase in energy consumption and a decrease in detection probability. However, $R_{s}$ declines compared to when $N_{\rm tx}=12$, indicating a higher blocklength for $N_{\rm tx}=16$. This occurs because common power coefficients are used to maintain a favorable beam pattern, and increasing the number of transmit APs may constrain power due to some APs being closer to the UEs. As a result, a higher blocklength is necessary to fulfill the reliability requirements. These findings underscore the importance of optimizing the number of transmit and receive APs.
The results of detection probability is depicted in Fig. 5(b). In general, a decrease in blocklength translates to a lower detection probability explaining the results for $N_{\rm tx}=12$. However, we observe that detection probability for $N_{\rm tx}=16$ still decreases although blocklength is higher compared to $N_{\rm tx}=12$. This is because the power coefficients required to satisfy the sensing SINR requirement might be lower with higher $N_{\rm tx}$, resulting in a more pronounced negative impact on detection probability. This negative impact outweighs the benefits of increased blocklength, leading to a lower detection probability.
The findings emphasize the importance of optimizing the number of transmit and receive APs to balance energy consumption, detection probability and sensing update rate. Higher numbers of transmit APs can enhance URLLC satisfaction and reduce blocklength resulting in higher sensing update rate, but must be carefully managed to avoid an unnecessary increase in energy consumption and detection probability degradation.
Figure 5: (a) E2E energy consumption, (b) detection probability, and (c) sensing update rate vs. the number of transmit ISAC APs for $N_{\rm rx}=1,2,4$ and E2E SeURLLC+ algorithm.
@@ VIII Conclusion
In this paper, we proposed a joint blocklength and power control algorithm for a downlink CF-mMIMO system with multi-static sensing and URLLC UEs. We formulated two non-convex optimization problems focused on minimizing transmission and E2E energy consumption, and conducted GOPS analysis for both communication and sensing tasks. Numerical results demonstrated the effectiveness of the E2E energy minimization algorithm, especially in scenarios without sensing capabilities. Integrating sensing reduces the optimal blocklength compared to URLLC-only scenarios due to the higher power demands of sensing, which leads to a lower optimal blocklength. Moreover, the E2E algorithms minimize total energy consumption by further reducing the blocklength. Additionally, we observed that higher sensing SINR thresholds significantly lower the sensing update rate without substantially improving detection probability. Our findings also indicated that increasing the number of sensing receive APs raises sensing-related processing energy, suggesting that deactivating some APs can help reduce E2E energy consumption in CF-mMIMO ISAC systems. While increasing the number of transmit APs can enhance URLLC satisfaction and sensing update rate by reducing blocklength, careful management is necessary to prevent excessive energy consumption and degradation in detection probability.
@@ Appendix A Proof of Theorem 1
Let us define a new optimization variable, denoted by $\overline{L}$, where $L-L_{p}\leq 1/\overline{L}$. Then, minimizing the objective function is equivalent to minimizing the following convex function (quadratic-over-linear function) $\frac{\|\bm{\rho}\|^{2}}{\overline{L}}$ Minimizing this function, at the optimal solution, holds that $\overline{L}=1/L_{d}=1/(L-L_{p})$. The reliability constraints in (51b) can be written in the form of
@@
according to (31), where $\overline{\mathsf{SINR}}^{\rm(dl)}_{i}$ is substituted by (28). To handle the non-convexity of the left-hand side in (61), we define a new variable $\chi_{i}$ and use fractional programming [44] to write the left-hand side as
@@
which can be written as a second-order cone (SOC) constraint in (52d). We then re-cast the constraint in (62) as
@@
$\bm{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\bm{\rho}\!\leq\!-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}.$ (65)
Finally, we define $\textbf{F}_{k}=\textrm{diag}\left(\sqrt{\mathbb{E}\left\{\|\textbf{w}_{0,k}\|^{2}\right\}},\ldots,\sqrt{\mathbb{E}\left\{\|\textbf{w}_{N_{\rm ue},k}\|^{2}\right\}}\right)$ and rewrite the per-AP power constraints in (51e) in SOC form in terms of $\bm{\rho}$ as $\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},$ , for $k=1,\ldots,N_{\rm tx}$.